OpenStudy (anonymous):

The width of the door is 4 feet. The door and wall form a right angle. C=(pi)(d), pi= 22/7 What distance does the edge of the door move when the door is opened all the way? A. 4 feet B. 5 3/7 feet C. 6 2/7 feet D. 7 feet E. None of the above

jimthompson5910 (jim_thompson5910):

imagine a circle that represents the door's path  with a radius of 4 feet

jimthompson5910 (jim_thompson5910):

what is the circumference of the full circle?

OpenStudy (anonymous):

C= 25.13?

jimthompson5910 (jim_thompson5910):

leave it in fraction form

jimthompson5910 (jim_thompson5910):

I'm assuming you used the formula C = 2*pi*r right?

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

so you should have C = 2*(22/7)*4 = ??? what is the result as a fraction?

OpenStudy (anonymous):

25.13 but as a fraction it would be 25 13/100?

jimthompson5910 (jim_thompson5910):

you should get 176/7 when you compute 2*(22/7)*4

jimthompson5910 (jim_thompson5910):

176/7 is the full circumference we just want 1/4 of it so multiply (176/7) by (1/4), then convert to a mixed number